OpenStudy (anonymous):

solve for x- x-1 over 6 = 1/3 choices are = 1. 1 2. -3 3. 3 4. 5

OpenStudy (anonymous):

x-1 1 --- = --- 6 3 Cross multiply. 3(x-1) = 6(1) 3(x-1) = 6 3x - 3 = 6 3x = 6 + 3 3x = 9 x=9/3 x=3

OpenStudy (anonymous):

That was a lot easier than it looked on my paper...